OpenStudy (anonymous):

Please Help With Factoring. 81m^4 + 72m^2n +16n^2

OpenStudy (anonymous):

Can Youu Explain Please.?

OpenStudy (anonymous):

\[(9m ^{2}+4n)^{2}\]

OpenStudy (anonymous):

Still Not Understanding...

OpenStudy (anonymous):

\[(a + b)^{2} = a ^{2} + 2ab + b ^{2}\]

OpenStudy (cwrw238):

compare your expression with the one that jeevan posted

OpenStudy (cwrw238):

a^2 = 81m^4 so a = 9m^2 now do the same with b^2 = 16n^2

OpenStudy (anonymous):

I Really Need An Explanation, Step To Step.

OpenStudy (cwrw238):

can you see what b = ?

OpenStudy (anonymous):

No.

OpenStudy (cwrw238):

having found b you substitute the values you ve found for a and b into (a + b)^2

OpenStudy (mertsj):